OpenStudy (anonymous):

Can you show the steps and answer to the following problem? Solve the system of equations by substitution. N_(Q) - 2ND = 11 3N_(Q) + N_(D) = -2

OpenStudy (anonymous):

Start by solving one of the equations for one variable in terms of the other variable. It doesn't matter which equation or variable--try to pick an easy one: Nq - 2Nd = 11 Nq = 11 + 2Nd Now substitute that value into the other equation and solve: 3Nq + Nd = -2 3(11 + 2Nd) + Nd = -2 33 + 6Nd + Nd = -2 7Nd = -35 Nd = -35/7 Now plug that into any of the previous equations to solve for the other variable: Nq = 11 + 2Nd Nq = 11 + 2(-35/7) Nq = 11 - 10 Nq = 1